OpenStudy (anonymous):

The legs of a right triangle measure 5√2 cm and 4√2 cm. Find the area.

OpenStudy (ivettef365):

do you have choices ?

OpenStudy (anonymous):

10√2 sq. cm 20 sq. cm 20√2 sq. cm

OpenStudy (mathstudent55):

In a right triangle, the right angle is formed by the legs. That means, for every right triangle, , one leg is the base and the other leg is the height. The area of a triangle is the product of the base and height divided by 2, so just multiply the lengths of the legs and divide by 2 to find the area.

OpenStudy (ivettef365):

A = 1/2 . b. h

OpenStudy (ivettef365):

what is 5\(\sqrt{2}\) x 4\(\sqrt{2}\)

OpenStudy (anonymous):

ok so 5√2 x 4√2 = area so 20 square root 4

OpenStudy (ivettef365):

ok, that is right, so what is square root of 4

OpenStudy (anonymous):

2

OpenStudy (ivettef365):

ok so you have 20 x 2 = ?

OpenStudy (ivettef365):

that is the base x height Area = 1/2 base x height, so what is half of it

OpenStudy (anonymous):

it's 20 square root 2 isn't it?

OpenStudy (mathstudent55):

\(A = \dfrac{bh}{2} \) \(A = \dfrac{5\sqrt{2} \times 4\sqrt{2}}{2} \) Ok so far?

OpenStudy (anonymous):

10 square root 2

OpenStudy (mathstudent55):

No. Just follow what I am doing from the response just above. We just plugged in the lengths of the legs in the formula. I did it above. Are you ok with that?

OpenStudy (anonymous):

yes

OpenStudy (ivettef365):

what is 20 x 2 equal to ????

OpenStudy (anonymous):

40

OpenStudy (mathstudent55):

\(A = \dfrac{20 \times \sqrt{2 \times 2}}{2}\) \(A = \dfrac{20 \times \sqrt{4}}{2}\) \(A = \dfrac{20 \times 2}{2}\) \(A = \dfrac{20 \times 2}{2}\) \(A = \dfrac{40 }{2}\) \(A = 20\) Answer: the area is \(20 ~cm^2\).

OpenStudy (anonymous):

ok then

OpenStudy (ivettef365):

the area is 1/2 of 40

OpenStudy (mathstudent55):

Keep this in mind: \(\sqrt{a} \times \sqrt{b} = \sqrt{ab} \) The way you use that rule is like this: What is \(\sqrt{2} \times \sqrt{8} = ~? \) \(\sqrt{2} \times \sqrt{8} = \sqrt{2 \times 8} = \sqrt{16} = 4 \)

OpenStudy (anonymous):

ok